OpenStudy (anonymous):

dentify the equation of the ellipse that passes through (4, 2) and has foci at (1, -1) and (1, 5)

OpenStudy (anonymous):

lol no one likes these conic section problems, but i bet we can do it

OpenStudy (tkhunny):

Didn't we already start on this one? Did you write the standard form, yet?

OpenStudy (anonymous):

your final answer will look like \[\frac{(x-h)^2}{b^2}+\frac{(y-k)^2}{a^2}=1\]

OpenStudy (anonymous):

oh you did this already?

OpenStudy (tkhunny):

Well, I was going to wait to discuss the placement of a and b, but you already fixed that. So, help2013, where does this leave us?

OpenStudy (anonymous):

thanks satellite73 and yes we did but im not sure